OpenStudy (anonymous):

the weights of bats in a zoo are normally distributed with a mean of 2.2 pounds and a standard deviation of 0.3 pounds. about what percent of the bats at the zoo weigh between 2.2 pounds and 2.5 pounds? 95% 34% 68% <--- is this the answer? 14%

OpenStudy (kirbykirby):

Should be half of 68%

OpenStudy (anonymous):

So it would be 34%? That's what. I got the first time but then I got 68% so I wasn't sure

OpenStudy (kirbykirby):

Yes When you standardize you get: \[P\left(\frac{2.2-2.2}{0.3}<\frac{X-2.2}{0.3}<\frac{2.5-2.2}{0.3}\right)=P(0<Z<1)\] So you're only looking from 0 to 1 standard deviation, not from -1 to 1 standard deviation (which is 68%)